OpenStudy (anonymous):

Which is equivalent to (3 + 2i)(3 - 2i)?

OpenStudy (tommo_lcfc):

(3+2i)(3-2i) = 6+4i^2 = 6 - 4 = 2 (as i^2 = -1)

OpenStudy (anonymous):

You almost had it tommo_lcfc. Niquita15, note that (3+2i)(3-2i) is the difference of squares. So you would have \[9-4i ^{2}\] Which becomes 9-[(4)(-1)] = 13

OpenStudy (tommo_lcfc):

What was I thinking? Silly mistakes! First 3*3 is not 6! Second, multiplying different signs together makes the answer negative! I am annoyed!

OpenStudy (tommo_lcfc):

VERY ANNOYED!!! I'm still smiling though :)

OpenStudy (anonymous):

-4 times -1 is a positive 4

OpenStudy (tommo_lcfc):

Yes but they are the same sign. -- and ++ = + +- and -+ = -

OpenStudy (anonymous):

Look at what I wrote down, and think about it.